OpenStudy (anonymous):

Simplify ( 2 square root 5 + 3 square root 7)^2. Show your work. Justify each step. Mom having a hard time understanding. Thank you for any help you can give

OpenStudy (phi):

I am not sure you can say you are simplifying... but you can "expand" it \[ (2 \sqrt{5} + 3\sqrt{7})^2 \] is the same as \[ (2 \sqrt{5} + 3\sqrt{7})(2 \sqrt{5} + 3\sqrt{7}) \] Do you know how to multiply this out (use FOIL if you know it) ?

OpenStudy (anonymous):

I do but I dont think I came up with the right steps for it